The following three defense stocks are to be combined into a price-weighted stock index in January...

  1. The following three defense stocks are to be combined into a price-weighted stock index in January 2016 (perhaps a portfolio manager believes these stocks are an appropriate benchmark for his or her performance): Suppose that Douglas McDonnell shareholders approve a 3-for-1 stock split on January 1, 2017 and that the company enacts the stock split on January 2, 2017.

    Price
    Shares
    (millions)
    1/1/16 1/1/17 1/1/18
    Douglas McDonnell 430 $ 108 $ 114 $ 41.08
    Dynamics General 535 38 34 48
    International Rockwell 210 67 56 70

    a. What is the new divisor for the index? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

    New Divisor:

    b. Calculate the rate of return on the index for the year of 2017 (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

    Rate of Return: %
Answer #1

a). Assume the divisor for 2018 be d

Index Value = [i=1(Pricei)] / n

a). Index Value(2017) = Index Value(2018)

[114 + 34 + 70] / 3 = [41.08 + 48 + 70] / d

218 / 3 = 159.08 / d

72.67 = 159.08 / d

d = 159.08 / 72.67 = 2.19

b). Index Value(2017) = [i=1(Pricei)] / n

= [114 + 34 + 70] / 3 = 218 / 3 = $72.67

Index Value(2018) = [i=1(Pricei)] / n

= [41.08 + 48 + 70] / 3 = 159.08 / 3 = $53.03

Return(2017) = [Index Value(2018) - Index Value(2017)] / Index Value(2017)

= [$53.03 - $72.67] / $72.67 = -$19.64 / $72.67 = -0.2703, or -27.03%